# Generative Video Semantic Communication via Multimodal Semantic Fusion with Large Model

### v1

@@ V Conclusion
In this work, we proposed the GVSC framework, a novel approach to video semantic communication using GenAI large models. This framework integrates video semantic extraction, source-channel transmission, and semantic encoding/decoding/generation. Video semantics is still an emerging concept, with varying interpretations across different tasks. In our paper, we adhere to the consistency of the transmitted semantic information. To address diverse transmission needs, we designed multimodal strategies. Compared with baselines like CLIP score, BERT score, PSNR, and SSIM, our schemes reduce transmission resources while achieving high-quality video reconstruction. We developed a weighted loss function to enhance the sketch transmission. Extensive simulations show that our schemes achieves a high semantic similarity at low CBR, demonstrating the feasibility of GenAI for efficient bandwidth-limited video transmission. While the proposed framework has no obvious advantages over conventional solutions in processing latency of video reconstruction, it holds significant potential under extremely low bandwidth and streaming media where instant interaction is not necessary. As generative model optimizations advance, we anticipate substantial enhancements in the real-time capabilities of our framework, promising a transformative impact on future communication systems.

### v2

@@ V Conclusion
In this work, we proposed the GVSC, a novel framework to video semantic communication using GenAI large models. It integrates video semantic extraction, source-channel transmission, and semantic encoding/decoding/generation. Video semantics is still an emerging concept, with varying interpretations across different tasks. We define video semantics as consistent semantic information during transmission. To handle diverse transmission needs, we designed multimodal strategies. Our method reduces transmission resources and achieves high CLIP, BERT, PSNR, and SSIM scores. We also design a weighted loss to improve sketch transmission. Extensive simulations show that GVSC works well at low CBR, demonstrating the feasibility of GenAI for efficient bandwidth-limited video transmission. Although it has higher latency than traditional methods, it is suitable for low-bandwidth and delay-tolerant scenarios. With future model optimizations, GVSC may support real-time applications and benefit future communication systems.